dedito22:

Manny has 48 feet of wood. He wants to use all of it to create a border around a garden. The equation 2l + 2w = 48 can be used to find the length and width of the garden, where l is the length and w is the width of the garden. If Manny makes the garden 15 feet long, how wide should the garden be?

Mercury:

"Manny makes the garden 15 feet long" so L = 15 therefore 2L + 2W = 48 plugging in L = 15 gives us 2(15) + 2W = 48 ^ try solving this for W let me know if you're having trouble or want to check your solution ^-^